OpenStudy (anonymous):

Simplify. 7x + 3(x - 2) - 4x + 8 A) 6x + 2 B) 6x + 14 C) 14x + 2 D) 14x + 14

OpenStudy (anonymous):

6x+2 = A

OpenStudy (anonymous):

thanks sooooo much..can u help me with more

OpenStudy (rane):

7x+ 3(x-2) - 4x 8 = 7x + 3x - 6 - 4x + 8 = 6x + 2

OpenStudy (anonymous):

Post it c:

OpenStudy (anonymous):

imma just post it under another one